51% of men consider themselves professional baseball fans. You randomly pick 10 men and ask each is he considers himself a professional baseball fan. Find the probability that the number who consider themselves baseball fans is
(a) exactly 5
(b) at least 6
(c) less than 4

Answer by Boreal(15235)

10 choose 5 * (0.51)^5(0.49^5)=0.2456
10 choose 6*(.51^6)(.49^4)=0.2130
10 choose 7*(.51^7)(.49^3)=0.1267
10 choose 8*(.51^8)(.49^3)=0.0494
10*(.51^9)(.49)=0.0114
(.51^10)=0.0012 TOTAL IS THE SUM OF THESE 5=0.4017
Probability
0: .49^10=0.0008
1: 10*(.51)*(.49^9)=0.0083
2: 45*(.51^2)(.49^8)=0.0389
3: 120*(.51^3)(.49^7)=0.1080 TOTAL IS 0.156